OpenStudy (anonymous):

1: given sin theta=4/5 with 90 dgreess less than or equal to thata less than or equal to 180 degress find exact value of : sin helppp pliss

OpenStudy (anonymous):

1: given sin theta=4/5 with 90 dgreess less than or equal to thata less than or equal to 180 degress find exact value of : sin

OpenStudy (anonymous):

\[\sin \theta=4/5, \theta=\sin^{-1} 4/5=53^{0}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i dont understand this

OpenStudy (anonymous):

first, you need theta and you have angles... you just have to find the inverse

OpenStudy (anonymous):

yes wat wre u ginna wrtite

OpenStudy (anonymous):

ok adn then